# MoVieDrive: Multi-Modal Multi-View Urban Scene Video Generation

### v1

@@ 3.4 Model Training and Inference
For model training, we employ the DDPM [10] noise scheduler and define the training objective $\mathcal{L}$ for multi-modal generation as:
@@
where for the $m$th modality, $\lambda_{m}$ is the weight for the training loss, $x_{0,m}$ is the groundtruth clean latent, $x_{t,m}$ is the noisy latent, $C$ is the condition, $\epsilon_{\theta,m}$ is the diffusion model, $t_{m}$ is the timestep. Besides, conditioning dropout is used as an augmentation strategy to improve model generalization and output diversity. During inference, random noises are sampled from the Gaussian distribution as the model inputs. We employ the DDIM [25] sampler to improve the efficiency of the reverse diffusion process and classifier-free guidance (CFG) [9] to balance output diversity and condition consistency.
Method Video Fidelity Scene Controllability Depth Quality Semantic Quality
@@
DriveDreamer 340.8 - - - - - -
DrivingDiffusion 332.0 - - - - - -
Drive-WM 122.7 - - - - - -
Panacea 139.0 - - - - - -
MagicDrive 217.9 11.5 18.3 84.4 0.255 72.8 23.5
DriveDreamer-2 55.7 - - - - - -
UniScene 70.5 - - - - - -
MaskGWM 65.4 - - - - - -
MagicDrive-V2 94.8 18.2 20.4 130.1 0.280 76.2 22.4
CogVideoX + SyntheOcc 60.4 15.9 28.2 51.6 0.124 298.9 32.4
Ours 46.8 22.7 35.8 25.3 0.110 36.2 37.5
Table 1: Quantitative comparison with the state-of-the-arts on nuScenes. Some results are cited from the original papers.
Figure 4: Qualitative comparison with the state-of-the-art methods. From left to right, we show results of back-right, back, back-left, front-left, front and front-right cameras and highlight some noticeable details.
@@ 4.1 Dataset and Experimental Setup
Following previous works [7, 6, 29], we conduct experiments on the challenging nuScenes [3] dataset. The nuScenes dataset is a real-world autonomous driving benchmark, containing 700 training videos and 150 validation videos. We use the official training and validation splits for experiments. For layout condition, we extract road maps and box maps following [29, 7] and prepare 3D occupancies following [14]. For reference conditions, we employ the first frames of all camera views following the common practice [29]. For text conditions, we utilize the camera intrinsic and extrinsic parameters as the camera prompt and prepare text prompts using CogVLM2-Caption [38]. For multi-modal training data preparation, we employ RGB videos from the official dataset and generate depth maps with Depth-Anything-V2 [36] and semantic maps with Mask2Former [4]. Note that other off-the-shelf models can also be used to prepare multi-modal training data. Since the goal of this work is not to surpass the state-of-the-art depth or semantic models but to demonstrate multi-modal multi-view scene generation in a unified model, we leave the integration of more state-of-the-art models for future work.
@@
For video fidelity evaluation, we use the commonly used Fréchet Video Distance (FVD) [28] as the metric. To evaluate the controllability, following [6], we use the pre-trained BEVFormer [16] to evaluate video-based 3D object detection in terms of mAP and video-based BEV segmentation in terms of mIoU. To evaluate the quality of the generated depth maps, we also use the Absolute Relative Error (AbsRel) [36]. Additionally, to evaluate the quality of generated semantic maps, we report mIoU for traffic-related classes, buildings, trees, and sky.
Implementation Details.
We build our model on CogVideoX(v1.1-2B) [38] and SyntheOcc [14]. We employ the pre-trained T5 text encoder [22] and initialize the 3D VAE and the temporal layers with the pre-trained weights from [38]. During training, the 3D VAE and the T5 text encoder are frozen, while the other model components are jointly trained. Our model is trained using the AdamW optimizer with a learning rate of $2e^{-4}$. By default, we use $6$ cameras and $49$ video frames, each with a resolution of $512{\times}256$.
Figure 5: Visualization of cross-modal consistency.
4.2 Comparison with the State-of-the-Arts
@@
We compare our approach with several state-of-the-art urban scene video generation methods, including MagicDrive [7], MagicDrive-V2 [6], DriveDreamer [29], DriveDreamer-2 [40], Drive-WM [31], Panacea [33], UniScene [13], MaskGWM [21], and DrivingDiffusion [15]. Besides, since we build our model on CogVideoX [38] and SyntheOcc [14], we use CogVideoX + SyntheOcc for multi-view scene video generation and consider it as our direct competitor. Moreover, since existing driving scene video generation methods lack the multi-modal generation capability, we use Depth-Anything-V2 and Mask2Former to estimate their depth maps and semantic maps. Note that single-view methods, such as Vista [8], are not our direct competitors because they cannot be directly used for multi-view video generation.
@@
Video fidelity measures the realism and temporal coherence of generated urban scene videos. We report FVD results in Tab. 1 and show that our approach achieves better results compared with the state-of-the-art methods. Specifically, our approach achieves FVD of 46.8, which is better than the state-of-the-art DriveDreamer, DriveDreamer-2, UniScene, MaskGWM, MagicDrive, etc. Besides, compared with CogVideoX + SyntheOcc, our approach improves performance by around 22%. Fig. 4 shows some qualitative comparison with the state-of-the-art methods. From Fig. 4, we can see that our approach can produce urban scene videos with high-fidelity details, e.g., vehicles and road structures.
@@
Scene controllability refers to the consistency between the conditioning inputs and the generated videos. We report mAP for 3D object detection and mIoU for BEV segmentation in Tab. 1. We can see that our approach achieves the best mAP of 22.7 for 3D object detection and the best mIoU of 35.8 for BEV segmentation, outperforming the state-of-the-art methods. This indicates that our approach can generate urban scene videos that are more consistent with the control conditions.
Result Analysis of Depth Quality.
In Tab. 1, we report FVD and AbsRel results to evaluate the quality of the generated depth map videos. Compared with the depth maps generated by Depth-Anything-V2, such as CogVideoX + SyntheOcc + Depth-Anything-V2, our approach achieves the better FVD of 25.3 and AbsRel of 0.110. This indicates that the multi-modal generation in our unified framework yields high-quality depth maps, even without using additional models for multi-modal generation.
Result Analysis of Semantic Quality.
In Tab. 1, we report FVD and mIoU results to evaluate the quality of generated semantic map videos. We can see that our approach achieves better results compared with the state-of-the-art methods, especially in terms of FVD which significantly outperforms the compared methods. This further verifies the effectiveness of our multi-modal generation approach.
@@
Table 2: Effectiveness analysis of our multi-modal generation. ‘✓’ means using our approach, ‘✗’ means using the off-the-shelf models.
4.3 Ablation Study
Effectiveness of Multi-Modal Generation.
In Tab. 2, we analyze the effectiveness of our multi-modal generation mechanism. We compare the performance of ours (RGB) + Depth-Anything-V2 + Mask2Former, ours (RGB + depth) + Mask2Former, and ours (RGB + depth + semantic). From Tab. 2, we can see that our unified approach (RGB + depth + semantic) achieves superior overall performance. Specifically, our approach achieves the best depth AbsRel and semantic mIoU and the comparable RGB FVD. This verifies that generating multiple modalities in a unified framework not only reduces the number of models, but also achieves better overall performance. We visualize some multi-modal generation results in Fig. 5. These results further show the cross-modal consistency of the multi-modal data generated from our approach.
Effectiveness of Diffusion Transformer Components.
In Tab. 3, we study the effectiveness of the proposed diffusion transformer model. We can see that using only temporal layer blocks (‘L1’) results in poor performance, while using temporal layer blocks plus modal-specific layer blocks (‘L1+L3’) brings better performance. With all components (‘L1+L2+L3’), our model yields the best performance. Besides, Fig. 6 shows that our approach can maintain cross-view consistency while our approach without using the multi-view spatiotemporal blocks yields poor cross-view consistency.
Effectiveness of Unified Layout Condition Encoder.
In Tab. 4, we analyze the effectiveness of our unified layout condition encoder. We can see that using the proposed unified layout condition encoder brings better performance compared with using pre-trained 3D VAE encoders. This can be attributed to the implicit condition embedding space alignment using the proposed module which effectively fuses the conditions for controllable scene generation.
@@ 4.3 Ablation Study
Table 3: Effectiveness analysis of the diffusion transformer components. ‘L1’: temporal layers, ‘L2’: multi-view spatiotemporal blocks, ‘L3’: modal-specific layers.
Figure 6: Visualization of cross-view consistency.
@@
Ours w/ the unified layout condition encoder 46.8
Ours w/o the unified layout condition encoder 53.4
Table 4: Effectiveness analysis of the unified layout condition encoder. ‘w/o the unified layout condition encoder’ means using pre-trained 3D VAE encoders.
@@ 5 Conclusion
This work presents a novel multi-modal multi-view urban scene video generation approach to autonomous driving. The key idea is to devise a new multi-modal multi-view diffusion transformer model and leverage diverse conditioning inputs to encode the controllable scene structure and content cues in a unified framework. Extensive experiments on the nuScenes dataset demonstrate the superiority of the proposed approach over the state-of-the-art methods.
Limitation and Future Work.
@@
While our approach has achieved superior multi-modal multi-view generation performance, there are still limitations. For example, how to effectively combine the proposed approach with a closed-loop autonomous driving simulator is worth further study. This potentially helps to comprehensively evaluate the safety and reliability of autonomous driving systems. In addition, how to derive LiDAR point clouds from the generated multi-modal data is an interesting research direction to further enhance our approach. Our future work aims to explore solutions to these problems to facilitate the deployment of our approach for real-world applications.

### v2

@@ 3.4 Model Training and Inference
For model training, we employ the DDPM [10] noise scheduler and define the training objective $\mathcal{L}$ as:
@@
where for the $m$th modality, $\lambda_{m}$ is the weight for the training loss, $x_{0,m}$ is the groundtruth clean latent, $x_{t,m}$ is the noisy latent, $C$ is the condition, $\epsilon_{\theta,m}$ is the diffusion model, $t_{m}$ is the timestep. Besides, conditioning dropout is used as an augmentation strategy to improve model generalization and output diversity. During inference, random noises are sampled from a Gaussian distribution as the inputs. We employ the DDIM [26] sampler to improve the efficiency of the reverse diffusion process and classifier-free guidance [12] to balance output diversity and condition consistency.
Table 1: Quantitative comparison with the state-of-the-arts on nuScenes. Some results are cited from the original papers. The “diverse conditions” adopted may vary across methods. Implemented with cross-view and cross-frame attention.
Method Venue Diverse Multi-modal? Fidelity Controllability Depth quality Semantic quality
@@
DriveDreamer [31] ECCV24 340.8 - - - -
DrivingDiffusion[16] ECCV24 332.0 - - - -
Drive-WM [33] CVPR24 122.7 - - - -
Panacea [35] CVPR24 139.0 - - - -
MagicDrive [8] ICLR24 236.2 9.7 15.6 0.255 23.5
DriveDreamer-2 [43] AAAI25 55.7 - - - -
UniScene [14] CVPR25 70.5 - - - -
MaskGWM [22] CVPR25 65.4 - - - -
MagicDrive-V2 [7] ICCV25 112.7 11.5 17.4 0.280 22.4
CogVideoX [40] + ICLR25+ 60.4 15.9 28.2 0.124 32.4
SyntheOcc [15] arXiv24
MoVieDrive (Ours) - 46.8 22.7 35.8 0.110 37.5
@@ 4.1 Dataset and Experimental Setup
Following the state-of-the-arts [8, 7, 31, 43], we mainly conduct experiments on the nuScenes [3] dataset, a real-world autonomous driving benchmark. On nuScenes, we use the official training and validation splits. For each training scene, we randomly select a sequence at each iteration. For the layout condition, we extract road maps and box maps following [31, 8] and prepare 3D occupancies following [15]. For context reference conditions, we employ the first frames of all camera views following the common practice [43]. For the text conditions, we utilize the camera intrinsic and extrinsic parameters as the camera prompt and prepare text prompts using CogVLM2-Caption [40]. For multi-modal data preparation, since no groundtruth multi-modal data are provided, following the common practice, we generate depth maps with Depth-Anything-V2 [38] and semantic maps with Mask2Former [5] for multi-modal training. In practice, when groundtruth multi-modal data are provided, our approach can be used to train a unified model for multi-modal multi-view scene generation. Here, we use these models because they are widely used for depth and semantic prediction in autonomous driving scene understanding, but other off-the-shelf models can also be used to prepare multi-modal data, and we leave the integration of more state-of-the-art models for future work. Additionally, we evaluate our approach on the Waymo dataset [28]. On Waymo, we use the official training and validation splits with the three front views for video generation.
Implementation Details.
We build our model on CogVideoX(v1.1-2B) [40] and SyntheOcc [15]. The 3D VAE encoder and decoder are from [40] and initialized with the pre-trained weights. For text condition encoding, we employ the pre-trained T-5 [23] as the frozen $E^{text}$ and construct $E^{cam}$ with an MLP-based encoder consists of two linear layers, where LayerNorm and SiLU are used between layers. For layout condition encoding, we use causal convolutional layers to construct causal resnet blocks following [40] and then construct the unified layout encoder with separate causal blocks for each condition and a shared causal resnet block for all conditions. For context reference condition encoding, we use the frozen 3D VAE as $E^{vae}$ but set the temporal dimension to one. We construct the multi-modal multi-view diffusion transformer with the architecture presented in Section 3.3, where the temporal layers and projection heads are initialized with the pre-trained weights from [40] and the other layers are randomly initialized. During training, the 3D VAE and the T5 text encoder are frozen, while the other model components are jointly trained. We train our model with the AdamW optimizer using a learning rate of $2e^{-4}$. By default, we use $6$ cameras and $49$ video frames, each with a resolution of $512{\times}256$.
4.2 Driving Scene Video Generation
@@
For video fidelity evaluation, we use the commonly used Fréchet Video Distance (FVD) [30] as the metric. To evaluate the controllability, following [7], we use the pre-trained BEVFormer [17] to evaluate video-based 3D object detection in terms of mAP and video-based BEV segmentation in terms of mIoU.
@@
We compare our approach with state-of-the-art urban scene video generation methods, including MagicDrive [8], MagicDrive-V2 [7], DriveDreamer [31], DriveDreamer-2 [43], Drive-WM [33], Panacea [35], UniScene [14], MaskGWM [22], and DrivingDiffusion [16]. Here, different methods propose to use different diverse conditions, so we mainly compare with the original methods instead of modifying them. Besides, we report results of CogVideoX [40] + SyntheOcc [15](with cross-view attention layers) for multi-view scene video generation and consider this model as our direct competitor. Note that single-view methods [9] are not our direct competitors since they cannot be directly used for multi-view video generation.
@@
Video fidelity measures the realism and temporal coherence of generated urban scene videos. We report FVD results in Tab. 1. We can see that our approach achieves better results compared with state-of-the-art methods. Specifically, our approach achieves FVD of 46.8, which is better than DriveDreamer, DriveDreamer-2, UniScene, MaskGWM, MagicDrive, etc. Besides, compared with CogVideoX + SyntheOcc, our approach improves performance by around 22%. Fig. 4 shows some qualitative comparison with state-of-the-art methods. From Fig. 4, we can see that our approach can produce urban scene videos with high-fidelity details, e.g., vehicles and road structures.
@@
Scene controllability refers to the consistency between the conditioning inputs and the generated videos. We report mAP for 3D object detection and mIoU for BEV segmentation in Tab. 1. We can see that our approach achieves the best mAP of 22.7 for 3D object detection and the best mIoU of 35.8 for BEV segmentation, outperforming state-of-the-art methods. This indicates that our approach can generate urban scene videos that are more consistent with the control conditions.
Figure 4: Quantitative comparison on nuScenes. From left to right, we show results of back-right, back, back-left, front-left, front and front-right cameras and highlight some noticeable details. Please check the supplementary material for an enlarged version.
Figure 5: Visualization of cross-modal consistency on nuScenes.
4.3 Multi-modal Multi-view Video Generation
The purpose of this experiment is to verify the superiority of using a unified model for multi-modal multi-view video generation compared to the common pipeline of using multiple models to generate different modalities. Our approach can simultaneously generate RGB videos, depth maps, and semantic maps using the proposed unified framework. Since existing driving scene video generation methods lack the multi-modal generation capability, we generate the RGB videos first and then use off-the-shelf models to estimate the depth maps and semantic maps.
Evaluation Metrics.
To evaluate the quality of the generated depth maps, we use the Absolute Relative Error (AbsRel) [38]. And to evaluate the quality of the generated semantic maps, we report the mean Intersection over Union (mIoU) for traffic-related classes, buildings, trees, and sky. The smaller AbsRel and the larger mIoU indicate the better quality of the generated depth and semantic maps.
Evaluation on Depth Quality and Semantic Quality.
In Tab. 1, we report the depth quality in terms of AbsRel and semantic quality in terms of mIoU. We can see that compared with the depth maps generated using a multi-stage pipeline, our approach achieves the best AbsRel of 0.110 and mIoU of 37.5. This shows that the multi-modal generation in our unified framework yields high-quality depth maps and semantic maps, even without using additional models for multi-modal generation. In addition, we visualize some multi-modal generation results in Fig. 5. These results further show the cross-modal consistency of the multi-modal data generated from our approach.
Table 2: Ablation study of multi-modal generation on nuScenes. ‘✓’: using our approach. ‘✗’: using the off-the-shelf models.
@@ 4.3 Multi-modal Multi-view Video Generation
Table 3: Effectiveness analysis of the diffusion transformer components on nuScenes. ‘L1’: temporal layers, ‘L2’: multi-view spatiotemporal blocks, ‘L3’: modal-specific layers.
@@
Figure 6: Visualization of cross-view consistency on nuScenes.
4.4 Ablation Study
Effectiveness of Multi-Modal Generation.
In Tab. 2, we analyze the effectiveness of our multi-modal components. We compare the performance of ours(RGB) + Depth-Anything-V2 + Mask2Former, ours(RGB + depth) + Mask2Former, and ours(RGB + depth + semantic). Tab. 2 shows that our unified approach (RGB + depth + semantic) achieves the best depth AbsRel and semantic mIoU and comparable RGB FVD. This verifies that generating multiple modalities in a unified framework can reduce the number of models and achieve better overall performance.
Effectiveness of Diffusion Transformer Components.
In Tab. 3, we study the effectiveness of the proposed diffusion transformer model. We can see that using only temporal layer blocks (‘L1’) results in poor performance, while using temporal layer blocks plus modal-specific layer blocks (‘L1+L3’) brings better performance. With all components (‘L1+L2+L3’), our model yields the best performance. Besides, Fig. 6 shows that our approach maintains good cross-view consistency, while removing the multi-view spatiotemporal blocks results in poor consistency.
Figure 7: Effectiveness of (a) unified layout condition encoder and (b) diverse conditioning inputs. ‘C1’: text conditions, ‘C2’: box and road map conditions, ‘C3’: occupancy conditions.
Figure 8: Long video generation without reference frames.
4.5 Further Analysis and Discussion
Effectiveness of Unified Layout Condition Encoder.
In Fig. 7(a), we analyze the effectiveness of our unified layout condition encoder. Here, ours without the unified encoder uses a pre-trained 3D VAE encoder for encoding. We can see that using the proposed unified layout condition encoder brings better performance compared with using pre-trained 3D VAE encoders. This can be attributed to the implicit condition embedding space alignment using the proposed module, which effectively fuses the conditions.
Effectiveness of Diverse Conditioning Inputs.
In Fig. 7(b), we examine the effectiveness of the diverse conditions. We can see that only using text prompt results in poor performance, while adding box and road map conditions brings significant performance improvement, which is aligned with existing work. Besides, the use of occupancy conditions also improves performance.
Long Video Generation without Reference Frames.
Although our approach is not specifically designed for long video generation, our model can be extended for long video generation, even without reference frames. This is benefited from the temporal compression mechanism in 3D VAE [40], the proposed diffusion transformer model and the diverse conditions. We visualize some results in Figs. 1(c) and 8. We can see that our approach can generate long videos with consistent scene layout and content. Nevertheless, we also observe the presence of noisy regions in far-distance areas.
Evaluation on the Waymo dataset.
Furthermore, we evaluate on the Waymo dataset. Tab. 4 shows that on Waymo, our approach achieves FVD of 61.6, outperforming the baseline (CogVideoX + Cross-view attention + Our conditions) and the state-of-the-art CogVideoX+SyntheOcc. As shown in Fig. 9, our approach can generate diverse driving scenes with high-fidelity on Waymo.
Table 4: Evaluation on Waymo.
@@ 4.5 Further Analysis and Discussion
Ours 61.6
CogVideoX + Cross-view attention + Our conditions 120.6
CogVideoX [40] + SyntheOcc [15] 82.3
Figure 9: Visualization of experimental results on Waymo.
Driving Scene Style Editing with Text Prompts.
Our approach can also be used to transfer scene style by editing text prompts. As shown in Figs. 1(b) and 9, by editing the scene text prompts, we can generate diverse driving scenes under different time and weather conditions.
@@ 5 Conclusion
This work presents a novel multi-modal multi-view urban scene video generation approach to autonomous driving. The key idea is to devise a new multi-modal multi-view diffusion transformer model and leverage diverse conditioning inputs to encode controllable scene structure and content cues in a unified framework. Extensive experiments on two autonomous driving datasets demonstrate the superiority of the proposed approach over state-of-the-art methods.
Supplementary Material
In this supplementary material, we discuss the limitation of this work in Section 6, present more discussions in Section 7, and present additional visualizations of multi-view multi-modal generation results in Section 8.
6 Limitation and Future Work
@@
While our approach has achieved superior multi-modal multi-view generation performance, there are still some limitations. First, how to effectively combine the proposed approach with a closed-loop autonomous driving simulator is worth further study. This potentially helps to comprehensively assess the safety and reliability of autonomous driving systems. Second, how to derive LiDAR point clouds from the generated multi-modal data is an interesting research direction to further enhance our approach. Moreover, as discussed in Section 4.5, although our approach can be extended for long video generation, there is still room for improvement in terms of long video quality. Our future work aims to address these problems to facilitate the deployment of our approach for real-world applications.
7 More Discussions
Is this a straightforward modification of CogVideoX with cross-view attention?
Our approach is not a straightforward modification of CogVideoX [40] for driving scene video generation. Although we use temporal layers and 3D VAE from CogVideoX, we significantly modify them with a unified multi-modal multi-view diffusion transformer model for autonomous driving scene generation and propose efficient diverse conditioning inputs encoding to improve scene video generation quality and controllability. A straightforward modification of CogVideoX is adding cross-view attention layers, specific heads, and our conditions to CogVideoX. However, as shown in Tab. 5, this yields worse performance compared with our design. Another simple variant is using the fixed CogVideoX and 3D VAE with learnable modal-specific heads but using neither our multi-view spatiotemporal blocks nor cross-view attention layers. From Tab. 5, we can see that this simple variant achieves significantly worse performance. In addition, combining CogVideoX with SyntheOcc [15] still yields worse results compared with our approach.
Table 5: Comparison with other model vairants on nuScenes.
[…]
Ours 46.8
CogVideoX + Cross-view attention + Our conditions 118.4
Fixed CogVideoX + Learnable head 364.1
CogVideoX + SyntheOcc 60.4
Table 6: Comparison with a modal-specific variant on nuScenes.
[…]
Ours 46.8
Modality-specific head based variant 109.4
Is this a simple extension of existing multi-modal synthesis methods?
Our approach is not a simple extension of existing multi-modal synthesis methods. Compared with existing methods, the innovation of this work lies in that we propose the first work that exploits diverse conditioning inputs and a unified multi-modal multi-view diffusion transformer model for multi-modal multi-view autonomous driving scene generation. This fills a gap left by existing works, facilitating urban scene understanding in autonomous driving. As discussed in Section 2, none of the existing diffusion-based multi-modal synthesis methods are specifically designed for multi-modal multi-view autonomous driving scene generation. It is not a trivial task to modify these methods for driving scene generation. For example, HyperHuman [18] is one of the recent works that explores diffusion-based multi-modal synthesis, but it is specifically devised for human generation so adapting this method to our problem requires modifying conditioning input encoders, diffusion transformer layers, training processes, etc. Similar situations applied to other existing methods. To alleviate this concern, we implement a model variant that employs the modality-specific heads [37], multi-view spatiotemporal blocks, CogVideoX, 3D VAE, and ours conditions. As shown in Tab. 6, this variant still yields worse performance compared with our approach. Although more careful modifications of existing multi-modal synthesis methods may improve their performance, from this result we know that adapting existing general methods to multi-modal multi-view driving scene video generation is not a trivial task. We leave more comprehensive studies for future work.
Figure 10: Multi-modal multi-view generation with DVIS++ and VDA for multi-modal data preparation.
Can this approach be used with other off-the-shelf models for multi-modal data preparation?
Our approach is orthogonal to the used off-the-shelf models. In practice, when groundtruth multi-modal data are provided, our approach can be used to train a unified model for multi-modal multi-view scene generation. Since no groundtruth multi-modal data are provided on the nuScenes and Waymo datasets, we use Depth-Anything-V2 and Mask2Former in this work. These models are widely used for depth and semantic prediction in autonomous driving scene understanding. It is worth noting that RGB video generation is key for this task. If the quality of generated RGB frames is poor, multi-modal prediction can hardly yield good performance. Thus, the collaboration of multi-modal multi-view generation is a highlight of our approach and Tab. 2 verifies that our unified approach can reduce the number of models used and achieve better overall performance for multi-modal multi-view urban scene synthesis. To alleviate this concern, we further conduct experiments with DVIS++ [42] for semantic maps preparation and VDA [4] for depth maps preparation. Experimental results show that our approach with DVIS++ and VDA yields FVD of 47.4, which is similar to the performance of our approach with Depth-Anything-V2 and Mask2Former (FVD of 46.8). In Fig. 10, we show some results of our approach with DVIS++ and VDA. We can observe that our approach can still generate high-quality multi-modal multi-view urban scene videos with cross-view consistency and cross-modal consistency. Note that colors of the multi-modal maps are depended on the predefined color palette of the off-the-shelf models and can be mapped to calculate the metrics.
8 More Visualization Results
In Figs. 11, 12, 13, 14, 15, 16, 17, 18 and 19, we show additional visualizations of our multi-view multi-modal generation results, including enlarged versions of some figures presented in the main paper.
Figure 11: Quantitative comparison on nuScenes (an enlarged version of Fig. 4).
Figure 12: Additional visualizations of our multi-view multi-modal generation results on nuScenes (arbitrarily selected frames covering different time).
Figure 13: Additional visualizations of our multi-view multi-modal generation results on nuScenes (arbitrarily selected frames covering different time).
Figure 14: Additional visualizations of our multi-view multi-modal generation results on nuScenes (arbitrarily selected frames covering different time).
Figure 15: Additional driving scene generation with diverse weather conditions on nuScenes.
Figure 16: Additional long video generation without reference frames on nuScenes.
Figure 17: Additional long video generation without reference frames on nuScenes.
Figure 18: Additional visualizations of experimental results on Waymo.
Figure 19: Additional visualizations of our multi-view multi-modal generation results on Waymo (arbitrarily selected frames covering different time).